OpenStudy (anonymous):

For the simple harmonic motion equation d=5sin((pi/4)t), what is the frequency? If necessary, use the slash (/) to denote a fraction.

OpenStudy (anonymous):

is frequency same as period? then it is (2pi)/(pi/4)

OpenStudy (anonymous):

that will actually be 8 period is 8 maybe frequency is 1/8?

OpenStudy (kropot72):

\[d=5\sin \frac{\pi}{4}t=5\sin 2\pi ft\] Therefore: \[2\pi f=\frac{\pi}{4}\] \[f=\frac{\pi}{4}\times \frac{1}{2\pi}=\frac{1}{8}Hz\]

OpenStudy (anonymous):

@kropot72 can you explain your first step please...?

OpenStudy (kropot72):

The equation in the question can be rewritten as\[d=5\sin 2\pi ft\] where 2pi f is called the angular frequency for a SHM. \[\frac{\pi}{4}=2\pi f\]